A Bibliography of Books by Past and Present University of Utah Mathematics Department

Members

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 December 2023 Version 1.27

Title word cross-reference

R^{3} × [0, T] [71].

1965 [96]. 1972 [94]. 1982 [85]. 1985 [89]. 1988 [90]. 1989 [82]. 1990 [88].

2000 [81]. 2002 [86].

60th [81, 87].

= [66].

Advanced [33, 34, 91, 96]. Advancement [90]. advances [91]. Algebra [11, 67, 52]. Algebraic [73, 81]. Algebras [72]. American [90]. AMP [57].

AMS [80, 88]. AMS-IMS-SIAM [80]. Analysis [19, 39, 40, 41, 75, 16]. Analytical [33]. annual [90]. Applications [42, 51, 53, 54, 94, 95, 96]. Applied [44]. Approximation [44]. Approximations [18]. Areas [56]. Army [95, 96]. Aspects [27, 28, 84]. Association [90]. Asymptotic [17, 95]. August [85, 91]. automating [63]. automatisieren [63]. aux [62].

Ben [66]. Bifurcation [51, 69]. Biological [53, 58, 91]. Biology [10, 35, 36, 92, 89, 90]. birthday [81, 87]. built [4]. Bursting [8]. Business [52].

C. [81]. Calculus [1, 2, 3, 11]. Case [92]. celebration [81]. Cell [10, 92]. Center [95, 96]. Central [93]. Chances [27]. Change [19]. Change-Point [19]. Characterization [24, 25]. Chemistry [96, 89]. Chern [67]. City [81, 94]. Classes [67]. Classic [61, 63]. Clemens [81]. Combinatorial [31]. Compact [5]. Complex [82, 73, 74]. Composite [14]. Composites [55, 56]. Computation [6, 59]. Computational [33]. Computer [21, 22, 23]. Concentration [17]. conducted [95, 96]. Conference [80, 54, 86, 85, 76, 94]. Connections [73]. Continuation [69]. Continuous [93]. control [16]. Convergence [24, 25]. course [60]. Cyclic [57].

d’Alembert [77]. Data [42]. David [87]. Definite [93]. Delay [94]. deviations [60]. Dictyostelium [57]. Differential [11, 20, 54, 88, 59, 94, 69, 16, 95]. Diffraction [79]. Diffusing [28]. Diffusive [29]. discoideum [57]. Dispersal [58]. Distributions [43]. Doctrine [27]. Domain [38]. Domains [12, 77]. Dynamics [1, 2, 3, 29, 90].

ecological [4]. Ecology [92]. ecosystems [4]. Eigenvalue [69]. Electrical [86]. Elliptic [69, 70]. Embeddings [68]. Empirical [17]. Engineering [33]. England [91]. environment [4]. Equadiff [85]. Equation [77]. Equations [11, 20, 54, 88, 59, 94, 69, 71, 16, 95]. ETOPIM [86]. Excitable [34, 90]. Experiment [30]. Experimental [91]. Extending [56]. Exterior [77].

festschrift [83]. Fields [9, 49]. Fluids [78]. Formation [53, 91]. Foundations [75]. Frequency [38]. FRG [85]. Frontiers [53]. Function [6]. Functional [42, 94]. Fundamental [5].

G [83]. Gambling [26, 27]. Games [26]. Genesis [8]. Geometrical [31]. Geometry [88, 73, 81, 82]. German [63]. Gibbs [60]. Global [51]. Gratings [79]. Groups [5, 73, 80, 87, 76]. Guide [66].

H [81]. Handbook [6]. held [82, 34, 85, 86, 88, 89, 90, 91]. Herb [81]. honor [81, 87].

Ihre [63]. IMS [80]. Inequalities [51]. Inference [42]. Inhomogeneous [86]. Institute [96]. Integrity [30]. Interfaces [29]. Internal [29]. International [86, 85]. Introduction [37, 38, 49, 62, 60].

Jiao [66]. jobs [63]. Joint [80]. Jr. [87]. July [86]. June [80].

Kähler [5]. Kernels [93]. Kingdom [34, 91]. Klassische [63]. Kurtz [83].

Lab [21, 22, 23]. Lake [81]. large [60]. Layers [29]. Learning [21, 22, 23]. Leeds [34]. Library [6]. Lie [80, 82, 73]. Life [1, 2, 3]. Limit [19, 43, 93]. Linear [11]. Linux [63]. Linux-Tasks [63]. Local [67].

Madison [95, 96]. Manifolds [5]. Maple [11]. Mappings [12]. March [81, 88]. Markov [24, 25, 83]. Materials [14, 32]. MathCW [6]. Mathematical [6, 14, 26, 28, 84, 36, 45, 46, 47, 48, 53, 92, 90]. Mathematical-Function [6]. Mathematics [11, 32, 33, 37, 38, 44, 53, 95, 96]. Matrix [39, 40, 41]. May [82, 89, 90, 95]. Measure [72]. measures [60]. Mechanics [59, 96]. Media [9, 34, 86, 90]. meeting [89, 90]. Methods [15, 33, 36, 69]. Millennium [13]. Minicourse [20]. Modeling [1, 2, 3, 38, 92]. Modelling [14]. Models [53, 58]. Multiparameter [49]. Multiplicities [67]. Multiscale [32]. Multivariable [11].

Nan [66]. NATO [34, 91]. Navier [71]. Nervous [8]. Neural [9]. Neurons [9, 37, 38]. Nevada [90]. Nonlinear [16, 34, 35, 88, 89, 69].

Obstacle [51]. October [96]. Operator [43]. Operator-Limit [43]. Optical [86]. Optimal [26]. Optimization [15]. Oscillations [35, 57, 89]. Other [56]. Oxford [91].

Park [94]. Partial [20, 88, 69]. Pattern [53, 91]. Period [12]. Perturbation [96]. Physiology [84, 45, 46, 47, 48, 92]. Play [26]. Point [19]. Population [36]. Portable [6]. Positive [93, 70]. powłoki [64]. Principles [44, 4]. Prize [13]. Probabilistic [27]. Probability [1, 2, 3, 18, 43, 50, 93]. Problems [13, 51, 69, 70]. Proceeding [54]. Proceedings [34, 85, 86, 88, 91, 89, 76, 95, 96, 82, 90]. Processes [10, 17, 24, 34, 49, 25, 83]. Products [93]. Programmierung [63]. Programming [6]. Programowanie [64]. Propagation [78]. Properties [86].

quantum [96]. questions [90].

Radar [7]. Random [49]. Reacting [28]. real [80]. reductive [80, 87, 76]. related [83]. Reliability [17]. Representation [80, 76]. Representations [87]. Research [80, 34, 91, 95, 96]. Rhythm [8].

Salt [81]. SAS [21, 22, 23]. Scattering [77, 79]. Science [56, 90]. Scientists [1, 2, 3]. Script [66]. Scripting [61, 63]. scripts [62]. Semilinear [70]. seminar [96]. September [34]. Session [88]. Several [73]. Shōkai [65]. Shell [61, 63, 62, 66]. Shell-Programmierung [63]. shieru [65]. SIAM [80]. Single [9]. Sixth [86]. skryptów [64]. Snowbird [80, 86]. Software [6]. Solutions [70, 95]. Some [90]. Sonar [7]. Sound [78]. Special [88]. States [95, 96]. Statistics [18]. Stochastic [10, 20, 54]. Stokes [71]. Stratified [78]. Structural [15, 30]. Studies [26, 92]. Study [66, 69]. sukuriputo [65]. Summer [80]. Sundance [82]. Suspensions [57]. Symposium [81, 90, 82, 95]. System [8]. Systems [28, 58].

Tasks [63]. Tensor [93]. Textbook [48]. Their [94, 95]. Theorems [19, 93]. Theoretical [96, 91]. Theory [17, 30, 43, 55, 56, 93, 77, 79, 80, 82, 76, 96]. Thomas [83]. Topics [14, 40, 83]. Topological [68, 69]. Topology [31]. Transformation [44]. Transport [86].

Unilateral [51]. United [34, 91, 95, 96]. University [95, 96, 81, 89, 76]. Unix [63]. Unix/Linux [63]. Unix/Linux-Tasks [63]. Urban [4]. USA [86]. Using [6]. UT [86]. Utah [81, 82, 89, 94, 80, 76].

Variables [73, 74]. Variational [15, 51]. Vector [11]. Vegas [90]. via [69]. Vogan [87].

Wave [34]. Waves [9]. weekend [81]. Weighted [18]. Wisconsin [95, 96]. Workshop [34, 91]. Würzburg [85].

Xi [66]. Xue [66].

Zhi [66].

References

Adler:1998:MDL

[1] Frederick R. Adler. Modeling the Dynamics of Life: Calculus and Probability for Life Scientists. Brooks/Cole Publishing Co., Pacific Grove, CA, USA, 1998.

Adler:2005:MDL

[2] Frederick R. Adler. Modeling the Dynamics of Life: Calculus and Probability for Life Scientists. Thomson Brooks/Cole, Belmont, CA, USA, second edition, 2005.

Adler:2013:MDL

[3] Frederick R. Adler. Modeling the Dynamics of Life: Calculus and Probability for Life Scientists. Brooks/Cole, Cengage Learning, Boston, MA, USA, third edition, 2013.

Adler:2013:UEE

[4] Frederick R. Adler and Colby J. Tanner. Urban ecosystems: ecological principles for the built environment. Cambridge University Press, Cambridge, UK, 2013.

Amoros:1996:FGC

[5] J. Amorós, M. Burger, K. Corlette, D. Kotschick, and D. Toledo. Fundamental Groups of Compact Kähler Manifolds, volume 44 of Mathematical Surveys and Monographs. American Mathematical Society, Providence, RI, USA, 1996.

Beebe:2017:MFC

[6] Nelson H. F. Beebe. The Mathematical-Function Computation Handbook: Programming Using the MathCW Portable Software Library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017.

Blahut:1992:RS

[7] Richard E. Blahut, Willard Miller, Jr., and Calvin H. Wilcox. Radar and Sonar, volume 32 of IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992.

Bressloff:2005:BGR

[8] Paul C. Bressloff and Stephen Coombes, editors. Bursting the Genesis of Rhythm in the Nervous System. World Scientific Publishing Company, Singapore, 2005.

Bressloff:2013:WNM

[9] Paul C. Bressloff. Waves in Neural Media: From Single Neurons to Neural Fields. Lecture Notes on Mathematical Modelling in the Life Sciences. Springer, Dordrecht, The Netherlands, 2013.

Bressloff:2014:SPC

[10] Paul C. Bressloff. Stochastic Processes in Cell Biology, volume 41 of Interdisciplinary Applied Mathematics. Springer International Publishing, Cham, Switzerland, 2014.

Carlson:1997:MMM

[11] James A. Carlson and Jennifer M. Johnson. Multivariable Mathematics with Maple: Linear Algebra, Vector Calculus and Differential Equations. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997.

Carlson:2003:PMP

[12] James Carlson, Stefan Müller-Stach, and Chris Peters. Period Mappings and Period Domains, volume 85 of Cambridge studies in advanced mathematics. Cambridge University Press, Cambridge, UK, 2003.

Carlson:2006:MPP

[13] James A. Carlson, Arthur Jaffe, and Andrew Wiles, editors. The Millennium Prize Problems. American Mathematical Society, Providence, RI, USA, 2006.

Cherkaev:1997:TMM

[14] Andrej Cherkaev and Robert Kohn, editors. Topics in the Mathematical Modelling of Composite Materials. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1997.

Cherkaev:2000:VMS

[15] Andrej Cherkaev. Variational Methods for Structural Optimization, volume 140 of Applied mathematical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000.

Clarke:1999:NAD

[16] Frank H. Clarke, Ronald J. Stern, and Gert Sabidussi, editors. Nonlinear analysis, differential equations, and control, volume 528 of NATO science series. Series C, Mathematical and physical sciences, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. Kluwer Academic Publishers Group. Proceedings of the NATO Advanced Study Institute and Séminaire de Mathemátiques Supérieures on Nonlinear Analysis, Differential Equations, and Control, Montréal, Canada, July 27–August 7, 1998.

Csorgo:1986:ATE

[17] Miklós Csörgő, Sándor Csörgő, and Lajos Horváth. An Asymptotic Theory for Empirical Reliability and Concentration Processes, volume 33 of Lecture Notes in Statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986.

Csorgo:1993:WAP

[18] Miklós Csörgő and Lajos Horváth. Weighted Approximations in Probability and Statistics. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1993.

Csorgo:1997:LTC

[19] Miklós Csörgő and Lajos Horváth. Limit Theorems in Change-Point Analysis. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1997.

Dalang:2009:MSP

[20] Robert Dalang, Davar Khoshnevisan, Carl Mueller, David Nualart, and Yimin Xiao. A Minicourse on Stochastic Partial Differential Equations, volume 1962 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009.

Elliott:1995:LSC

[21] Rebecca J. Elliott. Learning SAS in the Computer Lab. Duxbury Press, North Scituate, MA, USA, 1995.

Elliott:2000:LSC

[22] Rebecca J. Elliott. Learning SAS in the Computer Lab. Duxbury Press, North Scituate, MA, USA, second edition, 2000.

Elliott:2010:LSC

[23] Rebecca J. Elliott and Christopher H. Morrell. Learning SAS in the computer lab. Brooks/Cole Cengage Learning, Boston, MA, USA, third edition, 2010.

Ethier:1986:MPC

[24] Stewart N. Ethier and Thomas G. Kurtz. Markov Processes: Characterization and Convergence. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1986.

Ethier:2005:MPC

[25] Stewart N. Ethier and Thomas G. Kurtz. Markov processes: characterization and convergence. Wiley series in probability and statistics; Wiley-Interscience paperback series. Wiley-Interscience, Hoboken, NJ, USA, 2005.

Ethier:2007:OPM

[26] Stewart N. Ethier and William R. Eadington, editors. Optimal Play: Mathematical Studies of Games and Gambling. Institute for the Study of Gambling and Commercial Gaming, Reno, NV, USA, 2007.

Ethier:2010:DCP

[27] Stewart N. Ethier. The Doctrine of Chances: Probabilistic Aspects of Gambling. Probability and Its Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010.

Fife:1978:MAR

[28] Paul C. Fife. Mathematical Aspects of Reacting and Diffusing Systems, volume 28 of Lecture Notes in Biomathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978.

Fife:1998:DIL

[29] Paul C. Fife. Dynamics of Internal Layers and Diffusive Interfaces, volume 53 of CBMS-NSF Regional Conference Series in Applied Mathematics. SIAM Press, Philadelphia, PA, USA, 1998.

Folias:1989:SIT

[30] E. S. Folias, editor. Structural Integrity: Theory and Experiment. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989.

Glaser:1970:GCT

[31] Leslie C. Glaser. Geometrical Combinatorial Topology, volume 1 and 2. Van Nostrand Reinhold, New York, NY, USA, 1970 and 1972.

Golden:1998:MMM

[32] Kenneth M. Golden et al., editors. Mathematics of Multiscale Materials, volume 99 of The IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998.

Gustafson:1998:ACM

[33] Grant B. Gustafson and Calvin H. Wilcox. Analytical and Computational Methods of Advanced Engineering Mathematics, volume 28 of Texts in applied mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998.

Holden:1991:NWP

[34] Arun V. Holden, Mario Markus, and Hans G. Othmer, editors. Nonlinear Wave Processes in Excitable Media, Proceedings of a NATO Advanced Research Workshop on Nonlinear Wave Processes in Excitable Media, held September 11–15, 1989, in Leeds, United Kingdom, volume 244 of NATO ASI series. Series B, Physics. Plenum Press, New York, NY, USA; London, UK, 1991.

Hoppensteadt:1979:NOB

[35] F. C. Hoppensteadt, editor. Nonlinear Oscillations in Biology. Lectures in applied mathematics; v. 17. American Mathematical Society, Providence, RI, USA, 1979.

Hoppensteadt:1982:MMP

[36] F. C. Hoppensteadt. Mathematical Methods of Population Biology, volume 4 of Cambridge studies in mathematical biology. Cambridge University Press, Cambridge, UK, 1982.

Hoppensteadt:1986:IMN

[37] F. C. Hoppensteadt. An Introduction to the Mathematics of Neurons, volume 6 of Cambridge studies in mathematical biology. Cambridge University Press, Cambridge, UK, 1986.

Hoppensteadt:1997:IMN

[38] F. C. Hoppensteadt. An Introduction to the Mathematics of Neurons: Modeling in the Frequency Domain, volume 6 of Cambridge studies in mathematical biology. Cambridge University Press, Cambridge, UK, second edition, 1997.

Horn:1985:MA

[39] Roger Horn and Charles R. Johnson. Matrix Analysis. Cambridge University Press, Cambridge, UK, 1985.

Horn:1991:TMA

[40] Roger Horn and Charles R. Johnson. Topics in Matrix Analysis. Cambridge University Press, Cambridge, UK, 1991.

Horn:2012:MA

[41] Roger A. Horn and Charles R. (Charles Royal) Johnson. Matrix Analysis. Cambridge University Press, Cambridge, UK, second edition, 2012.

Horvath:2012:IFD

[42] Lajos Horváth and Piotr Kokoszka. Inference for Functional Data with Applications, volume 200 of Springer Series in Statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012.

Jurek:1993:OLD

[43] Zbigniew J. Jurek and J. David Mason. Operator-Limit Distributions in Probability Theory. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1993.

Keener:1988:PAM

[44] James P. Keener. Principles of Applied Mathematics: Transformation and Approximation. Addison-Wesley, Reading, MA, USA, 1988.

Keener:1998:MP

[45] James P. Keener and James Sneyd. Mathematical Physiology, volume 8 of Interdisciplinary Applied Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998.

Keener:1998:MPI

[46] James P. Keener and James Sneyd. Mathematical Physiology, volume 8 of Interdisciplinary Applied Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998.

Keener:2000:MP

[47] James P. Keener and James Sneyd. Mathematical Physiology. Perseus Books, Cambridge, MA, USA, 2000.

Keener:2005:MTP

[48] James P. Keener and James Sneyd. A Mathematical Textbook of Physiology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2005.

Khoshnevisan:2002:MPI

[49] Davar Khoshnevisan. Multiparameter Processes: An Introduction to Random Fields. Springer Monographs in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002.

Khoshnevisan:2007:P

[50] Davar Khoshnevisan. Probability, volume 80 of Graduate studies in mathematics. American Mathematical Society, Providence, RI, USA, 2007.

Le:1997:GBP

[51] Vy Khoi Le and Klaus Schmitt. Global Bifurcation Problems in Variational Inequalities: Applications to Obstacle and Unilateral Problems, volume 123 of Applied Mathematical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997.

MacArthur:2014:BA

[52] Kelly MacArthur. Business algebra. Kendall Hunt Publishing Company, Dubuque, IA, USA, 2014.

Maini:2001:MMB

[53] Philip K. Maini and Hans G. Othmer. Mathematical Models for Biological Pattern Formation: Frontiers in Applications of Mathematics, volume 121 of The IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001.

Mason:1977:PCS

[54] J. David Mason, editor. Proceeding of the Conference on Stochastic Differential Equations and Applications. Academic Press, New York, USA, 1977.

Milton:2002:TC

[55] Graeme Milton. The Theory of Composites, volume 6 of Cambridge monographs on applied and computational mathematics. Cambridge University Press, Cambridge, UK, 2002.

Milton:2016:ETC

[56] Graeme W. Milton, editor. Extending the Theory of Composites to Other Areas of Science. Milton–Patton Publishers, P.O. Box 581077, Salt Lake City, UT 85148, USA, 2016.

Munk:1989:CAO

[57] Peter B. Munk and Hans G. Othmer. Cyclic AMP Oscillations in Suspensions of Dictyostelium discoideum, volume 323:1215 of Philosophical transactions of the Royal society of London. Ser. B, Biological sciences. Royal Society of London, London, England, 1989.

Othmer:1987:MDB

[58] Hans G. Othmer, Stephen R. Dunbar, and W. Alt. Models of Dispersal in Biological Systems. Sonderforschungsbereich 123, Stochast. Math. Modelle. Universität Heidelberg, Heidelberg, Germany, 1987.

Palais:2009:DEM

[59] Richard S. Palais and Robert A. Palais. Differential Equations, Mechanics, and Computation, volume 51 of Student mathematical library; IAS/Park City mathematical subseries. American Mathematical Society, Providence, RI, USA, 2009.

Rassoul-Agha:2015:CLD

[60] Firas Rassoul-Agha and Timo O. Seppäläinen. A course on large deviations with an introduction to Gibbs measures, volume 162 of Graduate studies in mathematics. American Mathematical Society, Providence, RI, USA, 2015.

Robbins:2005:CSS

[61] Arnold Robbins and Nelson H. F. Beebe. Classic Shell Scripting. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. Also available in Chinese [66], French [62], German [63], Japanese [65], and Polish [64] translations.

Robbins:2005:ISS

[62] Arnold Robbins and Nelson H. F. Beebe. Introduction aux scripts shell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. French translation of [61] by Eric Jacoboni.

Robbins:2006:KSP

[63] Arnold Robbins and Nelson H. F. Beebe. Klassische Shell-Programmierung: [automatisieren Sie Ihre Unix/Linux-Tasks]. (German) [Classic Shell Scripting: [automating your Unix/Linux jobs]]. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. German translation of [61] by Kathrin Lichtenberg.

Robbins:2006:PSP

[64] Arnold Robbins and Nelson H. F. Beebe. Programowanie skryptów powłoki. Helion, Gliwice, Poland, 2006. Polish translation of [61] by Przemysław Szeremiota.

Robbins:2006:SSS

[65] Arnold Robbins and Nelson H. F. Beebe. Shōkai shieru sukuriputo. Orairī Japan, Tōkyō, Japan, 2006. Japanese translation of [61] by Aoi Hyūga.

Robbins:2008:SJB

[66] Arnold Robbins and Nelson H. F. Beebe. Shell Jiao Ben Xue Xi Zhi Nan = Shell Script Study Guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. Simplified Chinese translation of [61].

Roberts:1998:MCC

[67] Paul Roberts. Multiplicities and Chern Classes in Local Algebra. Cambridge University Press, Cambridge, UK, 1998.

Rushing:1973:TE

[68] T. Benny Rushing. Topological Embeddings. Academic Press, New York, USA, 1973.

Schmitt:1982:SEB

[69] Klaus Schmitt. A Study of Eigenvalue and Bifurcation Problems for Nonlinear Elliptic Partial Differential Equations via Topological Continuation Methods. Université Catholique de Louvain, Institut de Mathématique Pure et Appliquée, Louvain, Belgium, 1982.

Schmitt:1994:PSS

[70] Klaus Schmitt. Positive Solutions of Semilinear Elliptic Problems. University of Montréal Press, Montréal, Québec, Canada, 1994.

Stenger:2016:NSE

[71] Frank Stenger, Don Tucker, and Gerd Baumann. Navier–Stokes Equations on R^{3} × [0, T]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016.

Taylor:1973:MAC

[72] Joseph L. Taylor. Measure Algebras, volume 16 of CBMS Regional Conference Series in Mathematics. American Mathematical Society, Providence, RI, USA, 1973.

Taylor:2002:SCV

[73] Joseph L. Taylor. Several Complex Variables with Connections to Algebraic Geometry and Lie Groups, volume 46 of Graduate Studies in Mathematics. American Mathematical Society, Providence, RI, USA, 2002.

Taylor:2011:CV

[74] Joseph L. Taylor. Complex Variables, volume 16 of Pure and applied undergraduate texts. American Mathematical Society, Providence, RI, USA, 2011.

Taylor:2012:FA

[75] Joseph L. Taylor. Foundations of Analysis, volume 18 of Pure and applied undergraduate texts. American Mathematical Society, Providence, RI, USA, 2012.

Trombi:1983:RTR

[76] Peter C. Trombi, editor. Representation theory of reductive groups: proceedings of the University of Utah conference 1982, volume 40 of Progress in mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1983.

Wilcox:1975:STD

[77] Calvin H. Wilcox. Scattering Theory for the d’Alembert Equation in Exterior Domains, volume 442 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975.

Wilcox:1984:SPS

[78] Calvin H. Wilcox. Sound Propagation in Stratified Fluids, volume 50 of Applied mathematical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984.

Wilcox:1984:STD

[79] Calvin H. Wilcox. Scattering Theory for Diffraction Gratings, volume 46 of Applied mathematical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984.

Arthur:2008:RTR

[80] James Arthur, Wilfried Schmid, and Peter E. Trapa, editors. Representation theory of real reductive Lie groups: AMS-IMS-SIAM Joint Summer Research Conference, June 4–8, 2006, Snowbird, Utah, volume 472 of Contemporary mathematics, Providence, RI, USA, 2008. American Mathematical Society.

Bertram:2002:SHC

[81] Aaron Bertram, James A. Carlson, and Holger Kley, editors. Symposium in honor of C. H. Clemens: a weekend of algebraic geometry in celebration of Herb Clemens’s 60th birthday, University of Utah, Salt Lake City, March 10–12, 2000, volume 312 of Contemporary mathematics, Providence, RI, USA, 2002. American Mathematical Society.

Carlson:1991:CGL

[82] James A. Carlson, C. Herbert Clemens, and David R. Morrison, editors. Complex geometry and Lie theory, Proceedings of a symposium held at Sundance, Utah, May 26–30, 1989, volume 53 of Proceedings of symposia in pure mathematics, Providence, RI, USA, 1991. American Mathematical Society.

Ethier:2008:MPR

[83] Stewart N. Ethier, Jin Feng, Richard H. Stockbridge, and Thomas G. Kurtz, editors. Markov processes and related topics: a festschrift for Thomas G. Kurtz, volume 4 of Institute of Mathematical Statistics collections, Beachwood, OH, USA, 2008. Institute of Mathematical Statistics. A four-day conference held at the University of Wisconsin, Madison, July 10–13, 2006, in celebration of Tom Kurtz’s 65th birthday.

Hoppensteadt:1981:MAP

[84] F. C. Hoppensteadt, editor. Mathematical Aspects of Physiology, volume 19 of Lectures in applied mathematics, Providence, RI, USA, 1981. American Mathematical Society.

Knobloch:1983:EPI

[85] Hans Wilhelm Knobloch and Klaus Schmitt, editors. Equadiff 82: Proceedings of the international conference held in Würzburg, FRG, August 23–28, 1982, volume 1017 of Lecture Notes in Mathematics, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. Springer-Verlag.

Milton:2003:PSI

[86] G. W. Milton, K. M. Golden, D. Dobson, and Z. V. Vardeny, editors. Proceedings of the Sixth International Conference on Electrical Transport and Optical Properties of Inhomogeneous Media: ETOPIM 6, held in Snowbird, UT, USA, 15–19 July 2002, volume 338(1–4) of Physica B, Condensed Matter, Amsterdam, The Netherlands, 2003. North-Holland.

Nevins:2015:RRG

[87] Monica Nevins and Peter E. Trapa, editors. Representations of reductive groups: in honor of the 60th birthday of David A. Vogan, Jr., volume 312 of Progress in mathematics, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2015. Birkhäuser.

Oliker:1990:GNP

[88] Vladimir Oliker and Andrejs Treibergs, editors. Geometry and Nonlinear Partial Differential Equations, Proceedings of the AMS Special Session held March 23–24, 1990, volume 127 of Contemporary Mathematics, Providence, RI, USA, 1990. American Mathematical Society.

Othmer:1986:NOB

[89] Hans G. Othmer, editor. Nonlinear oscillations in biology and chemistry: proceedings of a meeting held at the University of Utah, May 9–11, 1985, volume 66 of Lecture notes in biomathematics, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. Springer-Verlag.

Othmer:1989:SMQ

[90] Hans G. Othmer, editor. Some mathematical questions in biology: the dynamics of excitable media. Proceedings of the 1988 Symposium on Some Mathematical Questions in Biology, held at the annual meeting of the American Association for the Advancement of Science in Las Vegas, Nevada, May 4, 1988, volume 21 of Lectures on mathematics in the life sciences, Providence, RI, USA, 1989. American Mathematical Society.

Othmer:1993:ETA

[91] Hans G. Othmer, Philip K. Maini, and James D. Murray, editors. Experimental and theoretical advances in biological pattern formation, Proceedings of a NATO Advanced Research Workshop on Biological Pattern Formation, held August 27–31, 1992, in Oxford, England, United Kingdom, volume 259 of NATO ASI series. Series A, Life sciences, New York, NY, USA; London, UK, 1993. Plenum Press.

Othmer:1997:CSM

[92] Hans G. Othmer, Frederick R. Adler, Mark A. Lewis, and John C. Dallon, editors. Case Studies in Mathematical Modeling—Ecology, Physiology, and Cell Biology. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997.

Parthasarathy:1972:PDK

[93] K. R. Parthasarathy and Klaus Schmidt. Positive Definite Kernels, Continuous Tensor Products, and Central Limit Theorems of Probability Theory, volume 272 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972.

Schmitt:1972:DFD

[94] Klaus Schmitt, editor. Delay and Functional Differential Equations and Their Applications: Conference on Delay and Functional Differential Equations and Their Applications, Park City, Utah, 1972, New York, USA, 1972. Academic Press.

Wilcox:1964:ASD

[95] Calvin H. Wilcox, editor. Asymptotic solutions of differential equations and their applications; proceedings of a symposium conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, May 4–6, 1964, volume 13 of Publication of the Mathematics Research Center, U.S. Army, University of Wisconsin, New York, NY, USA; London, UK; Sydney, Australia, 1964. John Wiley and Sons.

Wilcox:1966:PTA

[96] Calvin H. Wilcox, editor. Perturbation theory and its applications in quantum mechanics; proceedings of an advanced seminar conducted by the Mathematics Research Center, United States Army, and the Theoretical Chemistry Institute at the University of Wisconsin, Madison, October 4–6, 1965, New York, NY, USA; London, UK; Sydney, Australia, 1966. John Wiley and Sons.